### 05 May 2026 v2

# GPT-Image-2 in the Wild: A Twitter Dataset of Self-Reported AI-Generated Images from the First Week of Deployment

Abstract
The release of GPT-image-2 by OpenAI marks a watershed moment in AI-generated imagery: the boundary between photographic reality and synthetic content has never been more difficult to discern. As generative models reach unprecedented levels of photorealism, the research community faces a pressing need for real-world datasets to benchmark and fine-tune AI-generated image detection systems—yet no public dataset of GPT-image-2 images exists. We introduce the GPT-Image-2 Twitter Dataset, the first published dataset of GPT-image-2 generated images—sourced from publicly available Twitter/X posts in the immediate aftermath of the model’s April 21, 2026 release. Leveraging the Twitter API v2 recent search endpoint, we collected 27,662 image records within a six-day window. After applying a multi-stage curation pipeline—including rule-based media filtering, a multilingual text heuristic classifier spanning English, Japanese, and Chinese creation-language signals, and browser-automated Twitter badge verification—we curate a high-confidence set of 10,217 confirmed GPT-image-2 images. A key negative result is that platform-level provenance signals (C2PA content credentials) are systematically destroyed by Twitter’s CDN on upload, with implications for the broader field of AI image attribution on social media platforms.
Figure 1: A sample of 30 images from the GPT-Image-2 Twitter Dataset, illustrating the breadth of content: anime illustrations, photorealistic portraits, text-heavy infographics, fantasy scenes, product mockups, food, nature, and architecture.
1. Introduction
The release of GPT-image-2 by OpenAI on April 21, 2026, constitutes a significant inflection point in the trajectory of AI-generated imagery. The model demonstrates a degree of photorealism and multilingual text coherence widely reported to surpass contemporary alternatives, precipitating a rapid and geographically distributed wave of creative activity across social media platforms. Twitter/X emerged as a primary venue through which users disseminated their outputs, compared generative models, and documented creative workflows within hours of the model’s public availability.
Datasets of real-world AI-generated images sourced from social media are categorically distinct from laboratory-controlled collections in at least two respects. First, they reflect the empirical distribution of creative prompts, subjects, and aesthetic choices employed by non-expert users in practice—a distribution that diverges substantially from curated academic benchmarks such as DiffusionDB [13] or GenImage [18]. Second, they capture the sociotechnical context of AI-generated content: the communities that produce it, the languages in which it is annotated, and the engagement dynamics it elicits—all of which are absent from synthetically constructed benchmarks.
The decision to initiate collection immediately following the model’s launch was driven by a critical temporal constraint: the Twitter API v2 recent search endpoint restricts access to content from the preceding seven calendar days. Any delay would have permanently foreclosed access to the earliest and most novel phase of GPT-image-2 deployment.
This paper makes the following contributions. First, we present a collection methodology grounded in multilingual creation-language queries designed to maximise precision over recall—targeting exclusively those tweets in which the author explicitly attributes the image to GPT-image-2. Second, we introduce a multi-stage curation pipeline that assigns collected records to one of three classes—confirmed, rejected, or uncertain—on the basis of structured text heuristics. Third, we conduct two provenance validation experiments: a negative result showing that C2PA content credentials are stripped by Twitter’s CDN, and a browser-automation badge-check of the uncertain tweet pool. Finally, we characterise the dataset’s statistical and visual properties across four complementary analyses.
2. Related Work
2.1. AI-Generated Image Datasets
The rise of large-scale generative models has motivated a growing body of work on dataset curation. LAION-5B [10] is a broad-coverage dataset of image-text pairs scraped from the web, portions of which contain AI-generated images as an incidental byproduct. More targeted collections have been constructed around specific models: DiffusionDB [13] assembled 14 million Stable Diffusion images with user-written prompts scraped from a community Discord server, while GenImage [18] is a million-scale benchmark pairing images from eight generators (Midjourney, Stable Diffusion, GLIDE, and others) with real photographs for detection research. Closest to our own work, Zhang et al. [17] released GPT4o-Receipt, a dataset of GPT-4o-generated document images collected for forensic analysis—the first model-specific dataset targeting an OpenAI image generator, and a direct precursor to our GPT-image-2 collection. These controlled collections differ from our social-media-first methodology in that they use ground-truth provenance through API generation or controlled scraping from known-provenance sources.
Our work is most directly comparable to social-media collection pipelines where provenance is established through creator self-report rather than ground-truth API generation. The key methodological challenge shared across such efforts is that tweet text provides the only provenance signal, since platform transcoding destroys embedded metadata.
2.2. AI Image Detection, Attribution, and Social Media Deployment
A substantial research thread addresses the detection of AI-generated images and their attribution to specific generators. Early detection work [12] trained classifiers on GAN outputs; subsequent work [2] demonstrated that such classifiers generalise poorly across generators. A recent comprehensive benchmark [9] evaluated open-source detection models out-of-the-box on diverse generators, finding substantial performance degradation against newer model families not seen at training time—a gap our dataset is positioned to help address for GPT-image-2 specifically. Model fingerprinting has been explored as an alternative [16], exploiting subtle spectral patterns left by specific architectures. Multimodal LLMs have also been investigated as zero-shot deepfake detectors [8], with mixed results that underscore the difficulty of generalising beyond training distributions. The C2PA standard [1], adopted by Adobe, Microsoft, and OpenAI among others, offers a cryptographic approach to content provenance through embedded metadata. Our dataset provides an empirical evaluation of C2PA-based verification in the Twitter context and documents a fundamental platform-level obstacle to its deployment.
The practical limitations of these detection and attribution approaches become particularly salient when examined in real-world social media settings [7]. An analysis of AI-generated political images on X during the 2024 US election [3] found that a small fraction of “superspreader” accounts drive the majority of AI image circulation. A study of AI-generated posts in Reddit art communities [11] found that transparently self-reported AI posts account for fewer than 0.5% of all image-based posts—a figure that contextualises our own observation of the silent creator majority. AMMeBa [barrón2024ammeba] is a large-scale survey and dataset of media-based misinformation in-the-wild, tracking the rise of generative AI content in fact-checked claims through late 2023. Taken together, these works illuminate the gap between laboratory detection performance and real-world deployment; we contribute a dataset specifically designed to support downstream AI-generated image research and document the collection methodology for reproducibility.
3. Data Collection and Curation
3.1. Platform and API Setup
All data were collected using the Twitter API v2 recent search endpoint [15], which provides access to publicly visible tweets from the preceding seven calendar days; our collection ran over a six-day window (April 21–26, 2026). Collection was performed under a Basic (paid) access tier, which charges \$0.005 per tweet read and allows up to 100 results per pagination request. We used the has:images operator to restrict results to tweets with at least one photo attachment, and appended -is:retweet to all queries to exclude retweets and focus on original posts.
3.2. Query Design and Collection
6 Query Types hashtags, @mentions, model name Twitter API v2 27,662 records Stage 1 Media Filter Stage 2 Text Heuristics Text-Confirmed 5,467 images Uncertain $\sim$21,500 Rejected $\sim$780 Provenance Validation Playwright / Chromium Badge-Confirmed 4,750 images No badge / unevaluable 1,147-1{,}14726,515
Figure 2: End-to-end collection and curation pipeline (left to right). Stage 1 removes non-photo media and failed downloads; Stage 2 applies multilingual text heuristics yielding three classes. Uncertain tweets are further checked via Playwright/Chromium browser automation for Twitter’s “Made with AI” badge, yielding 4,750 additional badge-confirmed images.
Query design prioritised precision over recall: any tweet mentioning the model name in a discussion or news context would satisfy a naive keyword search, so all queries were constructed to target tweets where the author actively describes the act of creating an image. Figure 2 illustrates the full pipeline. We assembled six query types, all filtered with has:images -is:retweet, summarised in Table 1.
Table 1: Query types used in collection with image yield and confirmation rate.
Query type Images Conf.%
Hashtag (#GPTImage2, #GPTimage2) 1,314 94%
ChatGPT + “prompt” 1,098 48%
Chinese creation signals 408 41%
English creation language 13,508 44%
Japanese creation signals 2,183 28%
Broad anti-noise 8,974 9%
Total 14,154 30.9%
The hashtag query achieves the highest confirmation rate (94%), reflecting the deliberate, community-facing nature of this annotation practice. English creation-language queries combine broad creation verbs (made, created, generated) with prompt-sharing signals (prompt:) and produce the largest raw volume. Japanese signals target post-positional constructions with で生成 (“generated with”); Chinese signals target 提示词 (“prompt”) and 生成的 (“generated”). The broad anti-noise query captures tweets that mention the model without explicit creation credit and accordingly has the lowest confirmation rate (9%).
A key empirical finding is that the search-indexed tweet pool saturates within the collection window. The Twitter recent search API does not index every public tweet; our collection reflects the search-indexed pool rather than the complete universe of GPT-image-2 posts on the platform, meaning the seven-day API window—not financial constraints—is the binding limit on collection volume for rapidly deployed models. All counts in this paper are at the image level (a single tweet may contribute multiple images); records are deduplicated by (tweet_id, media_key) pair, with cross-tweet near-duplicate images also removed.
3.3. Curation Pipeline
After downloading all matched tweets and media, we applied a two-stage filter. Stage 1 is a deterministic media filter: records where media_type is not photo (video thumbnails, animated GIFs) are removed (344 records), along with 803 records with failed image downloads. Using has:images rather than has:media in the API query itself reduced video thumbnails from 13% (observed in a pilot) to 1.5%, yielding 26,515 photo records for classification.
Stage 2 applies a rule-based text classifier to each record, assigning one of three labels. Rules are evaluated in priority order: rejection fires before confirmation, so a tweet containing both comparison language and creation credit is treated conservatively.
Confirmed.
The tweet text contains explicit creation-language alongside a recognisable model name variant. English signals include “made with,” “created with,” “generated by,” and “prompt:”. Japanese signals include 生成 and 作ってみた (“tried making”). Chinese signals include 提示词 and 生成的. AI art hashtags (#AIart, #AIイラスト) serve as weak confirmation signals when present.
Rejected.
The tweet shows comparison language (“vs” a competing model), references three or more distinct AI tools, or uses release-announcement phrasing (“released,” “now available”).
Uncertain.
All remaining records—mentions of GPT-image-2 without explicit creation credit or clear rejection signal. This is the dominant class ($\approx$67% of filtered records).
3.4. Classification Results and Provenance Validation
Table 2: Stage 2 classification results (deduplicated unique image records). The broader uncertain pool checked via Playwright (${\sim}21{,}500$) includes pre-deduplication records from the full 26,515-record Stage 2 pass.
Total Confirmed Rejected Uncertain Conf.%
14,154 4,187 396 9,571 29.6%
Only 4,187 of 14,154 filtered images (29.6%) meet the confirmed threshold (Table 2). The remaining 70% are not noise: they are tweets that genuinely share GPT-image-2 content but whose authors do not use explicit creation language—silent creators who post with minimal captions, standalone hashtags, or purely descriptive titles. This mirrors broader findings in AI image sharing: prior work [11] found that transparently self-reported AI posts represent fewer than 0.5% of image-based posts in Reddit art communities.
Recovering silent creators requires image-level or platform-level provenance signals. C2PA content credentials [5, 1] proved infeasible: Twitter’s CDN strips all embedded metadata on upload, leaving every image as a bare JPEG with no EXIF, XMP, or C2PA markers. As an alternative, Twitter/X’s “Made with AI” badge [14] provides a client-rendered platform signal. We checked all ${\sim}21{,}500$ uncertain tweets using a Playwright/Chromium browser-automation pipeline (3–5 s between requests; 12 s render wait per page; login walls and timeouts excluded). Of those that loaded successfully, 53.7% carried the badge, providing independent platform-side corroboration. Badge absence is not treated as disconfirmation, as creators can opt out or post before retroactive labelling was applied.
4. Dataset Statistics
4.1. Language and Temporal Distribution
The dataset is strongly multilingual (Table 3): English leads at 40%, followed by Japanese (27%) and Chinese (21%), which may reflect GPT-image-2’s strong multilingual text-rendering capability as a particular draw for non-English communities. Figure 3 shows the daily confirmed-image count. Peak activity fell on April 22 (1,175 images)—one day after the US-timezone launch—as global audiences in non-US time zones caught up. Activity declined steadily, reaching 204 images by April 26 as the novelty wave subsided.
Table 3: Language distribution of confirmed images (lang field).
Language Images %
English (en) 4,117 40.3%
Japanese (ja) 3,355 32.8%
Chinese (zh) 1,959 19.2%
French (fr) 258 2.5%
Other 528 5.2%
Total 10,217 100%
Apr 21Apr 22Apr 23Apr 24Apr 25Apr 2605005001,0001{,}0005725721,1751{,}175911911723723597597204204Date (April 2026)Confirmed images
Figure 3: Daily count of confirmed GPT-image-2 images in the collection window. Peak activity occurs on April 22 (Day 2), reflecting global uptake following the US-timezone release on April 21.
4.2. Image Resolution and Aspect Ratio
[…]
5. Visual Content Analysis
To characterise the semantic and visual diversity of the dataset we apply four complementary analyses: (1) CLIP zero-shot subject-matter classification, (2) OCR-based text-legibility detection, (3) face detection with demographic estimation, and (4) CLIP embedding clustering. All analyses operate on the full 10,217-image confirmed dataset.
5.1. Subject Matter Taxonomy
Figure 4: Subject matter distribution across 10,217 images classified via CLIP zero-shot. Text-graphic content dominates; fantasy/surreal and photorealistic portraits follow.
We classify all 10,217 images into eight coarse content categories using CLIP ViT-L/14 zero-shot classification with category-describing text prompts. Figure 4 shows the distribution. Text-graphic content dominates (33.2%, n=3,387n=3{,}387), encompassing posters, infographics, and typographic illustrations—likely reflecting users probing GPT-image-2’s well-publicised multilingual text-rendering capability. Fantasy and surreal scenes constitute 24.9% (n=2,542n=2{,}542), while photorealistic portraits account for 14.2% (n=1,453n=1{,}453). Anime and illustrated styles represent 10.9% (n=1,109n=1{,}109), driven predominantly by Japanese-language creators. Product visualisations and UI mockups (9.8%, n=1,000n=1{,}000) signal professional and commercial use even within the first week.
Figure 5: Representative examples from the eight subject-matter categories classified by CLIP zero-shot.
5.2. Text Legibility
Running EasyOCR (ZH/EN) on the full image set reveals that 82.0% of confirmed images contain legible text (8,383 of 10,217), with a median of 29 detected text regions per text-bearing image (Figure 6). This high prevalence is consistent with both the text-graphic category above and GPT-image-2’s well-publicised ability to render coherent multilingual text.
Figure 6: Text presence across the 10,217 confirmed images. Over four-fifths (82.0%) contain machine-readable text, reflecting GPT-image-2’s strong text-rendering capability.
Figure 7: Example images with high text density ($\geq$20 OCR-detected regions), illustrating GPT-image-2’s multilingual text rendering across poster, infographic, and typographic compositions.
5.3. Face Detection and Demographics
InsightFace (buffalo_l, ONNX/CUDA) detects at least one face in 59.2% of images (6,053 of 10,217), with 22,583 faces in total across those images (Figure 12). Among detected faces, the model estimates 53.4% male (12,055) and 46.6% female (10,528), with a mean estimated age of 42.3 years (median 40). These statistics are provided to characterise dataset content only. Face detection carries well-documented biases for non-photorealistic imagery: given that anime and illustrated styles constitute 10.9% of the classified images, face counts, gender estimates, and age distributions may be inflated in that subset. Demographic statistics should not be interpreted as reflecting real individuals.
Figure 8: Example images with multiple detected faces, spanning photorealistic portraits, animated characters, and group compositions.
5.4. Semantic Clustering via CLIP
We embed all 10,217 images with CLIP ViT-L/14 [6], reduce to 2-D via UMAP [4] (cosine metric, $k=30$ neighbours, min-dist 0.05), and cluster with HDBSCAN (min cluster size 15). The procedure yields 137 distinct semantic clusters (Figure 9); 33.2% of images are assigned to noise, reflecting the genuine visual diversity of the collection. The largest clusters contain 603, 220, and 175 images respectively, each corresponding to visually coherent aesthetic families (e.g., anime group portraits, typographic posters, photorealistic headshots). The language-coloured projection (Figure 9, right) shows partial but not complete language segregation: Japanese and Chinese creators overlap substantially in the anime cluster regions, while English creators are more evenly distributed across clusters.
Figure 9: UMAP projection of CLIP ViT-L/14 embeddings coloured by HDBSCAN cluster assignment (left) and by tweet language (right). 137 clusters emerge from 10,217 images; 33.2% are classified as noise, reflecting genuine visual heterogeneity.
Figure 10: Representative images from the four largest CLIP semantic clusters. Each cluster captures a visually coherent aesthetic family: anime group scenes (C7), typographic posters (C10), illustrated characters (C42), and photorealistic portraits (C37).
6. Limitations
A material caveat for downstream use is that GPT-image-2 supports both text-to-image generation (prompting from scratch) and image-to-image editing (modifying an existing photo). Our confirmation signals—creation-language phrases (“made with,” “generated by,” “prompt:”) and Twitter’s “Made with AI” badge—do not distinguish between the two: creators use the same vocabulary for both workflows, and Twitter applies the badge identically. The dataset is therefore a mixture of fully-generated and AI-edited images, and we do not have a reliable signal to separate them at scale because creators rarely disclose which workflow they used. This has implications for downstream research: detectors trained on this dataset will learn the signatures of both regimes together; attribution to GPT-image-2 specifically is harder for edited images where most pixels may be unmodified; and the face, OCR, and demographic statistics in Section 5 include features that may have been carried over from input photos rather than synthesised.
7. Conclusion
We have presented the GPT-Image-2 Twitter Dataset: 10,217 confirmed GPT-image-2 images collected from Twitter/X within six days of the model’s public release, curated through a multilingual text-heuristic pipeline spanning English, Japanese, and Chinese creation-language signals, browser-automated Twitter badge verification, and model name variant matching. We document that C2PA content credentials are systematically stripped by Twitter’s CDN, and that Twitter’s native “Made with AI” badge (present on 53.7% of checked uncertain tweets) provides a complementary platform-level provenance signal. The collection and curation code is released at https://github.com/scamai/gpt-image-2-dataset; the dataset itself is released separately at https://www.scam.ai/en/research.
Appendix A Additional Figures
Figure 11: Aspect ratio distribution (left) and native-resolution breakdown (right) for all 10,217 confirmed images. Portrait is the plurality at 53.5%. “Other” resolutions reflect Twitter CDN resampling on upload.
Figure 12: Face count distribution (left), gender split (centre), and estimated age histogram (right) across 10,217 confirmed images. 59.2% of images contain at least one detected face (22,583 total faces).